OpenStudy (paige):

A function f is defined by f: x-->e^x-1, where x>0 (1) State the range of f. (2) Find an expression for f^-1 (3) State the domain of f^-1

OpenStudy (paige):

Can anyone help me with this please?

OpenStudy (anonymous):

e^x has a particular shape, where it is only valid when x>0 and y>0. so the problem is all set there e^x - 1 means that the graph has been translated 1 up in the y-direction this means the range of f is y>1 instead of y>0, because the graph has been moved up 1 in the y-direction i am not sure what part 2 and 3 mean

OpenStudy (paige):

I checked the marking scheme and it says the anser is > e–1 or > 0.37

OpenStudy (anonymous):

question is for \[e^x-1\] or \[e^{x-1}\]?

OpenStudy (anonymous):

will make a significant difference. part 2 is asking for the inverse function, but i cannot answer without knowing which one you mean

OpenStudy (anonymous):

\[f(x)=e^x-1\] for x > 0 the range will be \[(-1,\infty)\]

OpenStudy (anonymous):

to find inverse set \[x=e^y-1\] and solve for y via \[x+1=e^y\] \[y=\ln(x+1)\] \[f^{-1}(x)=\ln(x+1)\]

OpenStudy (paige):

No, the question is e^x-1 (e to the power of x-1) ;)